OpenStudy (sasogeek):

A certain number with 2 digits is divided by the sum of the digits, the quotient of the result is 6 and the remainder is 5. the difference, between the given number and the number formed by reversing the given number, is 18. find the 2 digits

OpenStudy (anonymous):

Need hints first ?

OpenStudy (sasogeek):

no hints, that's the question and all the hints are in there

OpenStudy (sasogeek):

actually i didn't mean no hints lol, that's all there is to the question

OpenStudy (anonymous):

so what you want? complete solution ?

OpenStudy (sasogeek):

yeah :) the 2 digits and how u got them :) the only hint i can think of outside the question is that it's a simultaneous equation with 2 variables... quite obvious. my teacher asked us this and he wanted an answer in 5 minutes. a couple of people got the answer and explained it to me later but they didn't meet the time requirement so basically we all failed. he said he's going to bring such questions in our next class as a test... i need to get conversant with it.

OpenStudy (anonymous):

Start with \( 10x+y \)

OpenStudy (sasogeek):

first number is in tens and second number is in ones... 10x + y... makes sense

OpenStudy (sasogeek):

10x+y)/(x+y) = 6+ ( 5/(x+y) ) .... right?

OpenStudy (sasogeek):

i forgot the first bracket

OpenStudy (anonymous):

Okay there are two equations that you need to form: 1) \(10x+y = 6(x+y)+5 \implies 4x -5y =5 \) 2) \( 10x+y-10y-x = 18 \implies 9x-9y=18 \) Now, all you have to do is solve this, 5 mints is more than enough :)

OpenStudy (sasogeek):

i see, i guess i need more practice :) i just got back from school though so i'll go get my stuff and get back here and post more questions so u can check on me to see if i'm doing what's right :) thanks btw :)

OpenStudy (anonymous):

Fain I help :)